OpenStudy (anonymous):

2x=-10 Show steps? Answers: -5

OpenStudy (precal):

divide both sides by 2

OpenStudy (anonymous):

2x/2 = -10/2 x = -5

OpenStudy (precal):

2x means 2 times x what is the opposite of times? division

OpenStudy (anonymous):

just divided by 2